# Think Smart, Act SMARL! Analyzing Probabilistic Logic Driven Safety in Multi-Agent Reinforcement Learning

### v1

@@ 4.4.1. Variation in Shield Strength
SIPPO agents under $\mathcal{T}_{strong}$ demonstrate even greater cooperation, focusing more on stag hunts and less on plant harvesting. Despite similar stag penalties as the $\mathcal{T}_{weak}$ agents, they progressively improve safety, implying that they learn better strategies. Their rewards are significantly higher (by an order of magnitude) than those of agents using $\mathcal{T}_{weak}$, which are already much greater than the unshielded agents’ rewards.
4.4.2. Partially Shielded Population
Analogous to the fully shielded Markov Stag-Hunt experiments (Section 4.4), Figure 7 shows the agents’ behavior over time, with reward and safety graphs. The baseline IPPO results are the same as in Section 4.4, and results from the fully shielded population are omitted for clarity. The metrics displayed are: (7(a)) plant harvests, (7(b)) successful stag hunts with another agent, and (7(c)) solo stag hunts. The other graphs show training (7(d)) and evaluation (7(e)) rewards, and mean safety for shielded agents under $\mathcal{T}_{strong}$ (7(f)).
@@ 4.4.2. Partially Shielded Population
[…]
5. Discussion
While we believe the research presented in this paper provides valuable insights into the application of PLS in SMARL, several gaps and extensions must be acknowledged: ($i$) The first is that existing methods for solving probabilistic logic programs quickly become computationally expensive as the state and action spaces grow, making them infeasible to be used in real-world applications. This also leads to time-consuming hyperparameter tuning. Future work should also consider PLS being applied to larger agent populations (more than two agents). ($ii$) Next, the use of a-priori constraints in PLS provides predefined safety knowledge, which is straightforward for simple environments but becomes complex in dynamic, real-world systems. This approach limits scalability, relying on expert-defined safety measures that are time-consuming to develop and may not adapt well to rapidly changing or unforeseen situations, suggesting a need to focus on enabling systems to autonomously learn and update these constraints. ($iii$) The logical complexity of the ProbLog program in PLS impacts performance and behavior, as seen in the Markov Stag-Hunt experiments where the stronger shield promoted more cooperation but at the cost of increased complexity due to more hand-designed rules. Future work should examine how varying shield complexity affects computational load and safety outcomes. ($vi$) In MARL systems, asymmetric or partial shielding was explored where only shielded agents update parameters with safety gradients, while unshielded agents focus on maximizing rewards, optionally sharing parameters. A weighted training approach may balance safety and reward optimization by adjusting shared parameter updates according to whether agents are shielded or not. ($v$) finally, while PLS in SMARL aims to enhance safety, its implementation must carefully consider ethical and societal impacts, as hand-designed a-priori constraints may introduce biases, under-specified definitions, or system vulnerabilities that current research has yet to fully address.
Supplementary Material: Appendices
Appendix A Environment Specifics
A.1. Repeated NFGs (Stag-Hunt)
@@
In these experiments both agents choose an action simultaneously and receive a reward determined by the payoff matrix. Each episode is a series of 25 repeated games. Thus the agents must learn to maximize the discounted reward attained over the whole episode (Axelrod, 1980). The exact values in the payoff matrix can be found in Figure 8. The three Nash equilibria for the Stag-Hunt are $(Stag,Stag)$, $(Hare,Hare)$ and a mixed Nash equilibrium where each agent plays Stag 60% of the time and Hare 40% of the time.
Figure 8. Payoff/reward matrix instance for Stag-Hunt.
A.2. Centipede Game

### v2

@@ 4.4.1 Variation in Shield Strength
IPPO 18.74$\pm$4.98 0.09$\pm$0.11 5.50$\pm$2.04 13.71$\pm$4.97 16.07$\pm$5.80
SIPPO ($\mathcal{T}_{weak}$) 18.73$\pm$4.26 12.80$\pm$7.41 14.25$\pm$3.57 68.48$\pm$35.19 79.80$\pm$38.33 0.96$\pm$0.01
SIPPO ($\mathcal{T}_{strong}$) 11.33$\pm$4.08 77.32$\pm$8.82 11.07$\pm$3.25 386.86$\pm$46.50 393.07$\pm$65.27 0.95$\pm$0.01
To isolate the effect of constraint strength on emergent cooperation, we compare two shields: $\mathcal{T}_{weak}$ and $\mathcal{T}_{strong}$. These were constructed such that $\mathcal{T}_{strong}\subset\mathcal{T}_{weak}$ in terms of their safety constraints. The former constrains agents to cooperate only when both are near a stag, while the latter more strictly promotes cooperation regardless of position. Both shields are applied in a 2-agent setting using the SIPPO algorithm, with unshielded IPPO as a baseline.
In this experiment, analyzing reward or safety alone does not fully capture cooperative behavior, so the key metrics displayed in Table 3 are: total plant harvests per episode, successful stag hunts with another agent, and unsuccessful attempts to hunt a stag alone. Additionally, we show the total return for each episode, and mean safety for shielded agents. As expected, IPPO agents avoid cooperation (thereby avoiding risk) and consistently harvest plants for a small but reliable reward. SIPPO agents using $\mathcal{T}_{weak}$ show moderate improvement, achieving occasional stag hunts while maintaining similar plant collection. Their increased cooperation also results in more penalties, as agents are willing to risk solo hunts.
Under $\mathcal{T}_{strong}$, cooperation improves dramatically: agents prioritize stag hunting and harvest fewer plants, yet earn more than five times the episodic reward compared to SIPPO with $\mathcal{T}_{weak}$. This suggests that stronger shields not only induce safer behavior but also lead to better coordination and long-term returns.
4.4.2 Scaling Cooperation with Population Size
We next examine how cooperation and safety scale with group size when all agents are shielded using $\mathcal{T}_{strong}$. We vary the number of agents $n\in\{2,3,4,5\}$ and compare six algorithms: three unshielded baselines (IPPO, CSPPO, ACSPPO) and three shielded variants (SIPPO, SCSPPO, SACSPPO). Metrics include total plants harvested, stags hunted, solo hunt penalties, mean episodic safety, and per-agent reward.
Figure 5 shows that unshielded agents (IPPO, CSPPO, ACSPPO) consistently avoid cooperative behavior, harvesting plants for low-risk, low-reward outcomes. Notably, ACSPPO – despite its fully shared architecture – performs worst overall, with near-zero safety and minimal rewards, underscoring that parameter sharing alone is insufficient to induce prosocial behavior.
In contrast, all shielded variants show increasing cooperation as population size grows. SACSPPO agents, in particular, demonstrate the strongest cooperative intent: they achieve the highest number of stag hunts, while also incurring more penalties, reflecting increased willingness to take coordination risks. Despite these penalties, SACSPPO yields the highest rewards and very high adherence to the shield’s definition of safety constraints. While per-agent rewards tend to decrease with larger populations, this is expected given that key environment parameters – such as the number of stags and plants, grid size, and episode length – remain fixed across all settings.
These results demonstrate that PLS scales robustly with population size, especially when paired with full parameter sharing. SACSPPO leverages shared gradients and consistent safety signals to generalize cooperation more effectively across agents, even in more complex multi-agent settings.
Figure 5: Results for the final 50 episodes of Markov Stag-Hunt with all agents shielded with $\mathcal{T}_{strong}$, for $n\in\{2,3,4,5\}$.
4.4.3 Partial Shielding and Parameter Sharing
Finally, we investigate how cooperation emerges in partially shielded populations. Fixing the population size at $n=5$, we vary the number of shielded agents $k\in\{0,1,\dots,5\}$ and compare SIPPO, SCSPPO, and SACSPPO. All shielded agents use the strong shield $\mathcal{T}_{strong}$, and unshielded agents remain unconstrained. This experiment tests how shielded behavior propagates through shared parameters in populations where only a subset of agents receive explicit safety constraints.
Figure 6 shows that cooperation, as measured by total stags hunted, increases monotonically with the number of shielded agents across all methods. SACSPPO shows the strongest effect: even with just two shielded agents, it significantly outperforms the other methods in cooperative behavior and safety adherence. As more agents are shielded, SACSPPO reaches a high level of cooperation, but also accumulates the most stag penalties – suggesting increased risk-taking and cooperative intent, even when coordination occasionally fails.
Reward and evaluation performance trend similarly. While all methods improve with greater shield coverage, SACSPPO consistently yields the highest returns, particularly at higher shielded ratios. This indicates that the safety-driven cooperative strategies learned via PLS are not only prosocial but also effective (reward-maximizing). In contrast, SIPPO and SCSPPO agents remain more conservative: they hunt fewer stags, incur fewer penalties, and achieve lower overall rewards.
Plant harvesting does not uniformly decrease with increasing shielded ratio across all methods. While SIPPO and SCSPPO show a modest decline – suggesting a shift from self-interested strategies to coordinated risk-taking – SACSPPO exhibits an unexpected increase in plant harvesting as more agents are shielded. Taken together with SACSPPO’s high shield adherence, this may indicate that stronger parameter sharing drives homogenization that improves reward maximization, encouraging agents to opportunistically exploit lower-risk strategies. Alternatively, the behavior could reflect an emergent division of labor, where agents already near the stag focus on hunting while others stabilize returns via plant harvesting. Distinguishing between these hypotheses would require further analysis of agent roles and policy diversity. Importantly, mean safety steadily increases with shield coverage in all methods, but SACSPPO achieves the highest safety levels – highlighting its superior ability to align unshielded agents with shielded behavior through shared learning.
Together, these results demonstrate that PLS can induce prosocial coordination even under partial coverage, and that shared parameters are a key conduit for propagating safety-aligned behavior to unshielded agents.
Figure 6: Results for the final 50 episodes of the Markov Stag-Hunt with $5$ agents, varying the number of shielded agents using the strong shield $\mathcal{T}_{strong}$.
5 Related Work
In the domain of safe (MA)RL, a few but significant strides have been made recently in developing frameworks that ensure safety during learning and execution [15]. Gu et al. [14], propose constrained Markov games and a constrained multi-agent policy gradient optimization algorithm (MACPO), however without the use of any formal language semantics. [35] use Mealy machines to approximate a world model to ensure safety, leading to some inherent symbolic interpretability. Other related concepts include dynamic shielding, for instance, ElSayed-Aly et al. [12] who introduce two frameworks that dynamically enforce safety constraints – our own method is similar to their factored shields. Banerjee et al. [3], Zhang et al. [40] and [38] are recent works that use the notion of (dynamic) model predictive shielding, which use ‘backup policies’ when an unsafe action is attempted by an agent. Though these are on par with SOTA results, these also do not provide the interpretability benefits of formal language semantics. Melcer et al. [20] explore decentralized shielding (similar to factored shields), i.e. allowing each agent to have its own shield, reducing computational overhead and enhancing scalability. This and ElSayed-Aly et al. [12] use LTL as formal language, though with a high knowledge requirement of the underlying MDP – this is in contrast to our use of a PL framework which requires only safety-related parts of the states. Notably, Subramanian et al. [29] integrate PL neural networks as the function approximator within agents to affect interpretable top-down control; however, our method differs significantly as the agents may use any function approximator, maximizing potential learning flexibility.
6 Discussion and Conclusion
In this article, we introduced Probabilistic Logic Shielded MARL, a novel set of DQN- and PPO-based methods that extend PLS to MARL. Several classes of games were analyzed and shown to benefit from SMARL in terms of safety and cooperation, and theoretical guarantees of these methods were presented.
@@ 6 Discussion and Conclusion
Supplementary Material for Analyzing Probabilistic Logic Shields for Multi-Agent Reinforcement Learning
The supplementary material is organized in the following way:
  • Appendix
    A: Theoretical proofs of propositions.
  • Appendix
    B: In-depth discussion of environments used, including implementation details, reward functions and settings.
  • Appendix
    C: Details on shields including their ProbLog source codes and process of construction
  • Appendix
    D: Single-agent experiment used to as an initial indication of PLTD’s safety and efficacy.
  • Appendix
    E: Table and explanation of all hyperparameters used.
  • Appendix
    F: Training/reward curves.
Appendix A Proofs
A.1 Proof of Proposition 3.2
Proposition A.1.
A PL-SMARL algorithm with agents $i\in N$ with respective policies $\pi_{i}^{+}$ and shields $\mathcal{T}_{i}$ has a joint shielded policy $\vec{\pi}^{+}$ at least as safe as their base joint policy $\vec{\pi}$, and strictly safer whenever any individual base policy $\pi_{j}$ violates its shield $\mathcal{T_{j}}$.
Proof.
Let $\Pi_{i}$ be the space of all policies for each agent $i\in N$. We then recall from Yang et al. [39]’s Proposition 1 that $P_{\pi^{+}_{i}}(\texttt{safe}|s)\geq P_{\pi_{i}}(\texttt{safe}|s)$, $\forall\pi_{i}\in\Pi_{i}\text{ and }\forall s\in S$. Thus, according to the safety specifications in $\mathcal{T}_{i}$, every $\pi_{i}^{+}$ is at least as safe as $\pi_{i}$, and thus $\vec{\pi}^{+}$ is at least as safe as $\vec{\pi}$.
If all base policies $\pi_{i}\in\vec{\pi}$ already satisfy their respective shields $\mathcal{T}_{i}$ in all reachable states (i.e., $P_{\pi_{i}}(safe|s)=1$ for all $s$), then shielding leaves the policy unchanged: $\pi_{i}^{+}=\pi_{i}$. Otherwise, if any $\pi_{j}\in\vec{\pi}$ does not abide by $\mathcal{T}_{j}$, by construction, the shielded policy will assign higher probability to safe actions in expectation, i.e. $P_{\pi^{+}_{j}}(\texttt{safe}|s)>P_{\pi_{j}}(\texttt{safe}|s)$, and the resulting joint policy $\vec{\pi}^{+}$ will be strictly safer than $\vec{\pi}$. ∎
A.2 Proof of Proposition 3.4
First, for completion, we prove a lemma that states that any shield with perfect information (binary information about the valuations of the sensor(.) predicates) always yield $P_{\mathcal{T}}(\text{safe}\mid s,a)\in\{0,1\}$. Then we define a more specific form of Proposition 3.4 and show that PLTD converges under perfect sensor information and regular TD-learning assumptions.
Lemma A.2 (Perfect-information shield yields binary safety).
Let $\mathcal{T}$ be a ProbLog shield that, at a given state $s$, receives action facts action($a$) whose probabilities are the policy $\pi(a\mid s)$, and sensor facts $\ell$) that are deterministic in $s$ (i.e. we have perfect sensor information), i.e., $P_{\mathcal{T}}\!\bigl{(}\texttt{sensor}(\ell)=\mathrm{true}\bigr{)}\in\{0,1\}$ and $\texttt{sensor}(\ell)=\mathrm{true}\iff\ell\ \text{actually holds in }s.$
Suppose $\mathcal{T}$’s knowledge base contains, for each ground action $a$, a clause of the schematic form
@@ Appendix A Proofs
where the literal $\ell_{a}$ may be one of:
  1. i.
    a state-dependent ‘hazard’ literal, or
  2. ii.
    an always-true dummy literal (i.e. unsafe_next :- action($a$).).
Together with the complementary rule
@@
the shield satisfies
$P_{\mathcal{T}}\!\bigl{(}\texttt{safe\_next}\mid\texttt{action}(a)\bigr{)}\;\in\;\{0,1\}\qquad\text{for every }a.$ (5)
Define the state-wise safe-action set
$A_{\mathrm{safe}}(s)\;=\;\bigl{\{}\,a\mid P_{\mathcal{T}}(\texttt{safe\_next}\mid\texttt{action}(a))=1\bigr{\}}.$ (6)
Then
$P_{\mathcal{T}}(\text{safe}\mid s,a)=\begin{cases}1,&a\in A_{\mathrm{safe}}(s);\\[4.0pt] 0,&a\notin A_{\mathrm{safe}}(s).\end{cases}$ (7)
Proof.
Condition the ground program on action($a$). All remaining random facts are the deterministic $\ell$) atoms, so the program is now fully deterministic. Hence the queried probability is $1$ if and only if the rule for unsafe_next fails.
Case (a). If $\ell_{a}$ is false in $s$ the rule body fails, so safe_next holds with certainty.
Case (b). If $\ell_{a}$ is always true, the rule body succeeds, so safe_next is false with certainty.
Thus the conditional probability is always 00 or $1$, and partitioning actions by the value $1$ yields the stated set $A_{\mathrm{safe}}(s)$. ∎
Proposition A.3 (PLTD Convergence under Perfect Safety Information).
Let $\mathcal{M}=\langle S,A,P,r,\gamma\rangle$ be a finite MDP with $\gamma<1$ and let $\mathcal{T}$ be a PL shield with perfect sensor information. Assume the agent
  1. i.
    [GLIE] follows a time-varying policy $\pi_{t}$ such that (a) $\pi_{t}(a\mid s)=0$ for all $a\notin A_{\text{safe}}(s)$, and (b) $\pi_{t}\xrightarrow{t\to\infty}\pi^{+}$ while every safe pair $(s,a)$ is selected infinitely often (e.g. $\epsilon_{t}$-greedy on $A_{\text{safe}}(s)$ with $\epsilon_{t}\!\downarrow 0$, $\sum_{t}\epsilon_{t}=\infty$).
  2. ii.
    stores a tabular action–value function $Q_{\theta}$, and
  3. iii.
    updates $\theta$ by minimising the PLTD loss (Eq. 3.3) with a stepsize sequence $\{\eta_{t}\}$ satisfying $\sum_{t}\eta_{t}=\infty$ and $\sum_{t}\eta_{t}^{2}<\infty$ (regular TD-learning assumption, Sutton and Barto [30]).
Then, with probability $1$, the parameter sequence $\{\theta_{t}\}$ converges to a fixed point $Q^{\star}_{\mathrm{safe}}$ that is optimal over the set of safe policies, i.e.
$Q^{\star}_{\mathrm{safe}}(s,a)=$ $\max_{\pi:\,\pi(a^{\prime}\!\mid s)=0\ \forall\,a^{\prime}\notin A_{\mathrm{safe}}(s)}$ (8)
$\mathbb{E}_{\pi}\!\Bigl{[}\textstyle\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}\,\bigm{|}\,s_{t}=s,\;a_{t}=a\Bigr{]}.$
Proof.
Fix a finite MDP $\mathcal{M}=\langle S,A,P,r,\gamma\rangle$ with $\gamma\!<\!1$ and a ProbLog shield $\mathcal{T}$ whose sensor atoms are deterministic in the current state $s$; by Lemma A.2 the shield therefore induces binary state–action safety probabilities. Throughout, let $A_{\text{safe}}(s)\subseteq A$ denote the set of $s$-safe actions singled out in the lemma.
1. The PLTD loss collapses to the ordinary TD loss.
For any state–action pair $(s,a)$,
$P_{\mathcal{T}}\bigl{(}\text{safe}\mid s,a\bigr{)}\;=\;\mathbf{1}\!\{a\in A_{\text{safe}}(s)\}.$ (9)
Consequently $P_{\pi^{+}}(\text{safe}\mid s)=1$ for every $s$ that can be encountered under the behaviour policy $\pi^{+}$ (assumption i). Thus the safety penalty $S_{P}=\log P_{\pi^{+}}(\text{safe}\mid s)$ in Eq. 3.3 vanishes identically and the PLTD objective reduces to
$\mathcal{L}^{Q^{+}}(\theta)\;=\;\mathbb{E}_{d\sim\mathcal{D}}\!\bigl{[}\!\bigl{(}r^{t}+\gamma\mathcal{X}-Q_{\theta}(s^{t},a^{t})\bigr{)}^{2}\bigr{]},$ (10)
namely the usual squared one-step TD error (with $\mathcal{X}$ equal to $\max_{a^{\prime}}Q_{\theta}(s^{t+1},a^{\prime})$ in the off-policy variant or $Q_{\theta}(s^{t+1},a^{t+1})$ in the on-policy variant).
2. Behaviour–target consistency.
The agent acts with $\pi^{+}$ and evaluates TD targets under $\pi^{+}$, so there is no off-policy mismatch; the TD recursion is on-policy in the SARSA setting and uses the same transition-probability matrix in the Q-learning setting.
3. Almost-sure convergence of the tabular TD iterates.
Because $S$, $A$ and therefore the post-shield state–action space $\{(s,a)\mid a\in A_{\text{safe}}(s)\}$ are finite, rewards are bounded, and the stepsizes $\{\eta_{t}\}$ satisfy $\sum_{t}\eta_{t}=\infty$ and $\sum_{t}\eta_{t}^{2}<\infty$ (assumption iii), the classical stochastic-approximation proofs of Watkins and Dayan [36] for Q-learning or Sutton and Barto [30] for SARSA apply verbatim. A minor notational change is that the maximisation (in the off-policy case) is taken only over safe actions, but the contraction argument is identical. Under the usual further condition that every safe state–action pair is visited infinitely often (assumption i) the parameter sequence $\{\theta_{t}\}$ converges with probability 1 to the unique fixed point $Q^{\pi^{+}}$ of the Bellman operator corresponding to $\pi^{+}$.
4. Optimality of the limit among safe policies.
By construction $\pi^{+}$ places zero probability on unsafe actions and is greedy with respect to $Q^{\pi^{+}}$ over each safe-action set $A_{\text{safe}}(s)$:
$\pi^{+}(a\mid s)>0\;\Longrightarrow\;a\in\operatornamewithlimits{argmax}_{a^{\prime}\in A_{\text{safe}}(s)}Q^{\pi^{+}}(s,a^{\prime}).$ (11)
Standard policy-evaluation/-improvement arguments (e.g. Sutton and Barto [30]) now imply that $Q^{\pi^{+}}$ coincides with the optimal action-value function over the restricted policy class $\Pi_{\text{safe}}\;=\;\{\pi\mid\pi(a^{\prime}\mid s)=0\text{ whenever }a^{\prime}\notin A_{\text{safe}}(s)\}.$ We denote this limit by $Q^{\star}_{\text{safe}}$ and $\forall(s,a)$ obtain
$Q^{\star}_{\text{safe}}(s,a)\;=\;\max_{\pi\in\Pi_{\text{safe}}}\mathbb{E}_{\pi}\Bigl{[}\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}\,\bigm{|}\,s_{t}=s,\;a_{t}=a\Bigr{]},$ (12)
5. Putting it together.
Steps 1–4 establish that $Q_{\theta}$ converges to $Q^{\star}_{\text{safe}}$ as claimed.
Remark A.4.
As stated in the main paper, for tabular and linear function-approximation settings, Lemma A.2 together with the projected TD results of Tsitsiklis and Van Roy [33] yields convergence of PLTD. For deep neural Q-networks, global convergence is currently not known; in that regime we can adopt the usual stabilisation tricks (target network, replay buffer, Double-DQN). Empirically, we can show that PLTD remains stable across tasks, but a complete theoretical treatment is left for future work.
A.3 Proof of Proposition 3.5
Proposition A.5.
For any game there exists a shield $\mathcal{T}$ such that the resulting shielded agent deterministically selects a single action $a\in\mathcal{A}$ for each state $s\in S$ if the shielded policy $\pi^{+}$ computes all other actions $a^{\prime}\in\mathcal{A}\setminus\{a\}$ as perfectly unsafe.
Proof.
Assume a game $\mathcal{E}$ that has actions $\mathcal{A}$ for a given agent and $|\mathcal{A}|=n_{\mathcal{A}}$. This agent has a policy $\pi(s)=\langle a_{0},\vec{a}_{-a_{0}}\rangle\in\mathbb{R}^{n_{\mathcal{A}}}_{\geq 0}$, where $\vec{a}_{-a_{0}}$ is the tuple of action values except $a_{0}$, such that $\sum\pi(s)=1$. We recall from Definition … that for any agent with a policy $\pi$, a safe policy $\pi^{+}$ is computed as:
$\pi^{+}(a\mid s)=P_{\pi}(a\mid s,\mathtt{safe})=\frac{P(\mathtt{safe}\mid s,a)}{P(\mathtt{safe}|s)}\pi(a\mid s)$
We now construct a shield $\mathcal{T}$ in ProbLog. Let action(a0) be a probabilistic fact in $\mathcal{T}$ that represents $\pi(a_{0}|s)$ and with constraints $\mathcal{KB}$ as follows:
1action(0)::action(a0);
2... % optional: other actions
3
4% optional: sensor inputs
5
6% safety specification
7unsafe_next :- action(X), X\=a0.
8safe_next :- \+unsafe_next.
Note that the shield is state independent. Using the semantics of PL, we see that $P(\texttt{safe}|s,a^{\prime})=0$ for $a^{\prime}\neq a_{0}$, i.e. there is no state in which the safety of any other action apart from $a_{0}$ at all safe. Now, we can compute $\pi^{+}(a_{0}|s)$ and determine that $P(\texttt{safe}|s)=1-\sum\vec{a}_{-a_{0}}=a_{0}=\pi(a_{0}|s)$, leaving us with $\pi^{+}(a_{0}|s)=P_{\pi}(\texttt{safe}|s,a_{0})=1$. Thus, we get a safe deterministic policy (pure strategy) $\pi^{+}(s)=\langle 1,0,...\rangle$, sampling from which, the agent will always play $a_{0}$.∎
Appendix B Environment Specifics
B.1 Repeated NFGs (Stag-Hunt)
@@
In these experiments both agents choose an action simultaneously and receive a reward determined by the payoff matrix. With reference to the generic payoff matrix in 7, any game with a payoff matrix with the following structure of payoffs is considered an instance of the Stag-Hunt: $a=b>d=e\geq g=h>c=f$. Each episode is a series of 25 repeated games. Thus the agents must learn to maximize the discounted reward attained over the whole episode [2]. The three Nash equilibria for the Stag-Hunt game instance as in Figure 2 are $(Stag,Stag)$, $(Hare,Hare)$ and a mixed Nash equilibrium where each agent plays Stag 60% of the time and Hare 40% of the time.
Figure 7: Generic payoff matrix
B.2 Centipede Game